HelpMePlz:

megan invests $1,200 each year in an ira for 12 years in an account that earned 5% compounded annually. at the end if 12 years she stopped making payments to the account but continued to invest her acchnukated amount at 5% compounded annually for the next 11 years

Champion:

So A. 1200 for 12 years, then B. continues to accumulate for 11 years, right?

Champion:

Part A (annual payments) Future value (after 12 years) FV= A(1+R+R^2+...R^(n-1) ) =AR^n/(R-1) =AR^n/i A=annual payment i=annual interest rate, compounded yearly R=1+i (Here A=$1200, i=0.05, R=1.05, n=12) Part B (initial deposit, compound interest) Future value (after 11 years) = FV* R^n Here FV=result from part A R=1.05 n=11 (years)